Another kind of double salt of the sulfate anion is one with two kinds of cations: K

and Mg. The percent by mass of each component in schonite is given below:

% K = 19.42%

% Mg = 6.036%

% S = 15.93%

% H= 3.003%

%O = 55.62%

Determine the molecular formula of this double salt.
Answer #1

Moles of K = mass%/molar mass = 19.42/39 = 0.498

Moles of Mg = 6.036/24 = 0.25

Moles of S = 15.93/32 = 0.498

Moles of H = 3.003/1 = 3.003

Moles of O = 55.62/16 = 3.47

Molar ratio = K : Mg : S : H : O

= 0.498 : 0.25 : 0.498 : 3.003 : 3.47

= 2 : 1 : 2 : 12 : 14

Hence, molecular formula is K 2MgS2O14H12

Thus, on rearranging we get K 2Mg(SO 4)2.6H2O

Please give feedback!